# MAPLE: Enhancing Review Generation with Multi-Aspect Prompt LEarning in Explainable Recommendation

### v1

@@ A.1 Additional Details for Multi-Aspect Review Segmentation
For the aspect-tuple extractor, we use a T5-based fine-tuned information-extraction model uie-large-en  [^8] https://huggingface.co/luyaojie/uie-large-en(Lu et al., 2022) on SemEval datasets Pontiki et al. (2014, 2015, 2016). For the zero-shot aspect-category classifier for the aspect terms, we use facebook/bart-large-mnli released by Meta. We formulate the classification task as multiple Natural-Language Inference problems: for one aspect term, we we ask the question "{aspect-term}. This example is about {category}." for each aspect category, and choose the argmax-ed class out of their entailment logits as the final aspect category of this aspect term.
@@ A.2 Additional Dataset Details
For Yelp23, we downloaded all the reviews from the online Yelp hosted dataset website  [^9] https://www.yelp.com/dataset. Since the raw data contains a large number of reviews, we follow the guidelines of Li et al. (2020). That is, we recursively prune users and items with less than 20 interactions, and split the datasets randomly into training, validation and testing at the ratio of 8:1:1 for 5 times, while ensuring a warm-start scenario. Note that we split in terms of user-item pairs, i.e., all of the associated $(u,i,c_{1}),...,(u,i,c_{k})$ (k is the number of categories mined from the ground-truth review) are guaranteed to appear in one of the training, validation or testing. For Yelp19, we utilize the data as processed by Li et al. (2020). Our sentiment analysis pipeline is applied to both Yelp19 and Yelp23. For Yelp23, we retain text spans illustrated in 1 as $E_{u,i,c}$. Since Yelp19 consists of only fragments of full reviews–which are not available to us–we treat the entire text associated with each user-item pair as $E_{u,i,c}$.
@@ A.3 Aspect Category Inventory
Yelp23 African Cuisine, American Cuisine, Asian Cuisine, Barbecue and Steakhouses, Breakfast and Cafes, Burmese and Mongolian Cuisine, Chinese Cuisine, Comfort Food and Diners, European Cuisine, Food Court and Stands, Gastropubs and Modern European, "Gluten-Free, Vegan, Vegetarian", Greek and Mediterranean Cuisine, Halal and Kosher, Hot Pot, Japanese and Sushi, Korean Cuisine, Latin American Cuisine, Live/Raw Food and Salad, Middle Eastern Cuisine, Seafood, South Asian Cuisine, Southeast Asian Cuisine, Tapas Bars, ambience, location, miscellaneous; service
@@ B.1 Automatic Evaluation Metrics
    where $\hat{E}_{u,i}$ is the generated sentence for a user-item pair, $f_{u,i}$ is the set of features extracted from the ground-truth review, and $\delta(x)=1$ if $x$ is true, $\delta(x)=0$ otherwise.
For informativeness (iFMR) and some of the aspect-wise explainability metrics (iFCR, FCR), with an eye to encourage more fine-grained aspect terms and penalize the overly-generic terms, we filter out the keywords that are too short (less than 4 characters), too general or noisy (see below list of dummy_words).
@@ B.2 Additional Details for Baseline Models
  • PETER (Li et al., 2021): An un-pretrained transformer-based model that integrates user and item embeddings into the generation process.
  • PEPLER (Li et al., 2023b): A model leveraging a pre-trained GPT-2 with prompt tuning for generating explanations. We specifically use the PEPLER-MF variant, known for its superior text quality. Among these models, PEPLER’s architecture closely resembles ours, making it a crucial point of comparison.
@@
  • PRAG (Xie et al., 2023): A transformer-based model integrates user and item embeddings along with review history. It employs a personalized attention mechanism to generate a latent query, which is optimizes in an auto-encoder fashion against the embedding of the ground-truth review.
@@ C.1 Implementation Details
MAPLE is trained on the training set, with hyper-parameters fine-tuned on the validation set. Performance evaluation is conducted on a sampled subset of 10,000 pairs from the testing set for efficiency. For review-generation models and MAPLE, results are averaged over 5 data splits, while PRAG is evaluated only on the first split due to its extensive training time. MAPLE is optimized using the AdamW optimizer (Loshchilov and Hutter, 2019), with a learning rate of 0.001 and a batch size of 196. The maximum sequence length is limited to 20 Byte Pair Encoding (BPE) (Sennrich et al., 2016) tokens. The training process is divided into two stages as discussed in Section 3.6: stage 1 runs for up to 30 epochs with the tolerance times set to 5 (monitoring only the text loss); stage 2 lasts for a maximum of 20 epochs with the tolerance times set to 2 (monitoring only the aspect-recommendation loss). We use different training datasets for each stage. In the first stage, each user-item pair is associated with various aspect categories $c_{i}$ from the training set. In the second stage, focused on training the recommendation task, we trim the dataset so that each user-item pair appears only once. The aspect regularization coefficient $\alpha$ is set to 0.01. The embedding dimension for user ID, item ID and aspect is set to 768. For the aspect-recommendation component, we use a MLP with two hidden layers of 256 and 128 dimensions. Regarding the Distribution Balance loss, we set the negative-tolerance coefficient $\lambda$ to 2.0 and class bias $\nu_{i}$ to 0.05 for each class. During inference, we apply greedy decoding for all models. For the personalized retriever of PRAG, we employ its item-marginalization variant and train with user, item ID dimension of 768; the rest follow the default settings. For all the other baseline models, we adopt the official implementation as well as the default hyper-parameter settings.
C.2 Computational Resources and Costs
Experiments are run on a single NVIDIA GeForce RTX 3090 (24 GB) GPU. In the MAPLE +RAG variant in the case studies D, we run the retrieval by first encoding all of the reviews with a sentence transformer and then computing the similarity scores within a subset of review pool. This action requires around 60 GB of CPU RAM. For all +RAG case studies, we use the GPT-4 Achiam et al. (2023) model hosted on OpenAI ChatGPT website 101010urlhttps://chatgpt.com/ on June 10, 2024, with default setting and a cleared history.
@@ C.2 Computational Resources and Costs
Figure 5: Prompt template for the reader model. In terms of the personal query, for MAPLE+RAG, we use the MAPLE generated explanation; for PRAG+RAG, we use the translated and filtered keyword set.
@@ Appendix D Qualitative Case Studies
To address the issues of under-representations of user role, we select several examples under the same item for case study. Since the personalized retriever in Xie et al. (2023) produces a latent query for a user-item pair, we train the embedding interpreter as Xie et al. (2023) proposes to translate the query to a set of keywords: We input the latent query into the embedding interpreter and list the output keywords. In implementation, it generates 5 keywords and then keeps only those appearing in the retrieved reviews, therefore the displayed keywords may be less than five. In the enhanced +RAG model, we treat the prefix model as the retriever and utilize GPT-4 as the reader. We augment only MAPLE and PRAG with this setup. For MAPLE, we encode its explanation with all-mpnet-base-v2  [^11] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 and then treat them as queries to fetch the top-10 most similar reviews from the user and item history review pools, respectively. See 6 for illustration. For PRAG personalized retriever, we use its proposed item-marginalization variant (subtracting the mean of latent queries generated for the same item to highlight distinctiveness for each item) to retrieve pertinent reviews from the user and item pool.
D.1 Case Study 1: Item Feature Preciseness
It can be observe that in Table 7, PRAG’s retriever often produces queries focused narrowly on service speed, which lacks the breadth to capture the restaurant’s primary attractions. Consequently, the PRAG+RAG mostly centers the discussions on service and wait times. There are some more cases shows that PRAG has tendency to prioritize service aspects. Conversely, MAPLE excels in identifying detailed features such as ice cream flavors (e.g., salted caramel, chocolate chips, strawberry), styles (e.g., twist) and its complement (e.g., the waffle cone), offering a variety of explanations tailored to different user preferences. In contrast, other review-generation models generally produce generic terms like "ice-cream" or "toppings"; the best is probably PETER’s "soft-served" in recommending user xC-q_yh0XwcjRLimkS3RNg.
@@ D.2 Case Study 2: Personalization
In exploring the personalization capabilities of recommendation systems, we focus on how effectively restaurant features are remembered and presented in user-specific contexts, i.e., how item IDs impact the generated explanations. A pertinent question arises: How about the impacts of user IDs? We assume an imaginary a baseline approach, which simply samples reviews from an item’s review history. Could it possibly be as good as MAPLE or the other review-generation models? Or let’s put it this way, what are the differences between it or the other review-generation models? We observe that user signals primarily influence the style and tone of the generated sentences. For instance, in Table 8, the user has a review history predominantly in French. Learning that the past reviews are primary French, the models generate sentences also in French. This level of personalization would be challenging to achieve with this random sampling baseline from the item’s review history. Could such a baseline possibly rival in terms of personalization of MAPLE or other user ID-incorporated review-generation models? Moreover, MAPLE demonstrates an ability to retain and reflect personal details. For example, in Table 9, the user frequently mentions "Phoenix" in the review history, suggesting they might reside in the City of Phoenix. MAPLE captures this detail and incorporates it into the generated explanation, highlighting the user’s local preferences. This ability to integrate personal context into recommendations underscores the model’s strength in tailoring content to individual users. These observations lead to the conclusion that the random baseline could not replicate the nuanced personalization achieved by models like MAPLE, which possess sophisticated mechanisms for personalizing generated sentences, adapting not only to users’ linguistic preferences but also to subtle personal details, thereby enhancing the relevance and effectiveness of personalized recommendations.
@@
Table 7: Case studies on item id N_2yEZ41g9zDW_gWArFiHw (Sweet Jesus, an ice-cream shop) in Yelp19. There are total 35 reviews in the 10,000 test samples, 4 reviews are selected for demonstration. The recommended aspect terms mentioned in the explanations are boldfaced.

### v2

@@ A.1 Additional Details for Multi-Aspect Review Segmentation
For the sentiment-analysis model, we use a T5-based fine-tuned information-extraction model uie-large-en  [^9] https://huggingface.co/luyaojie/uie-large-en(Lu et al., 2022) on SemEval datasets Pontiki et al. (2014, 2015, 2016). For the zero-shot aspect-category classifier, we use the off-the-shelf facebook/bart-large-mnli released by Meta. We formulate the classification task as multiple Natural-Language Inference problems: for one aspect term, we ask the question "{aspect-term}. For each aspect category, we use the prompt "This example is about {category}.", and select the category with the highest entailment logit as the final aspect category for the given aspect term. [^10] https://huggingface.co/facebook/bart-large-mnli In this setup, we use T5 for precise span-level extraction, and BART for reliable category-level classification. We find this modular design to be more effective than relying on a single multi-task model.
@@ A.2 Additional Dataset Details
For Yelp23, we downloaded all the reviews from the online Yelp-hosted dataset website  [^11] https://www.yelp.com/dataset. Since the raw data contains a large number of reviews, we follow the guidelines of Li et al. (2020). That is, we recursively prune users and items with fewer than 20 interactions, and split the datasets randomly into training, validation, and testing at the ratio of 8:1:1 5 times, while ensuring a warm-start scenario. Note that we split in terms of user-item pairs, i.e., all of the associated $(u,i,c_{1}),...,(u,i,c_{k})$ (k is the number of categories mined from the ground-truth review) are guaranteed to appear in one of the training, validation, or testing stages. For Yelp19, we utilize the data as processed by Li et al. (2020). Our sentiment analysis pipeline is applied to both Yelp19 and Yelp23. For Yelp23, we retain text spans illustrated in 2 as $E_{u,i,c}$. Since Yelp19 consists of only fragments of full reviews–which are not available to us–we treat the entire text associated with each user-item pair as $E_{u,i,c}$.
@@ A.3 Aspect Category Inventory
Yelp23 African Cuisine, American Cuisine, Asian Cuisine, Barbecue and Steakhouses, Breakfast and Cafes, Burmese and Mongolian Cuisine, Chinese Cuisine, Comfort Food and Diners, European Cuisine, Food Court and Stands, Gastropubs and Modern European, "Gluten-Free, Vegan, Vegetarian", Greek and Mediterranean Cuisine, Halal and Kosher, Hot Pot, Japanese, and Sushi, Korean Cuisine, Latin American Cuisine, Live/Raw Food and Salad, Middle Eastern Cuisine, Seafood, South Asian Cuisine, Southeast Asian Cuisine, Tapas Bars, ambiance, location, miscellaneous; service
@@ B.1 Automatic Evaluation Metrics
    where $\hat{E}_{u,i}$ is the generated sentence for a user-item pair, $f_{u,i}$ is the set of features extracted from the ground truth review, and $\delta(x)=1$ if $x$ is true, $\delta(x)=0$ otherwise.
For informativeness (iFMR) and some of the aspect-wise explainability metrics (iFCR, FCR), with an eye to encouraging more fine-grained aspect terms and penalizing the overly generic terms, we filter out the keywords that are too short (less than 4 characters), too general or noisy (see below list of dummy_words).
@@ B.2 Additional Details for Baseline Models
  • PETER (Li et al., 2021): An unpretrained transformer-based model that integrates user and item embeddings into the generation process.
  • PEPLER (Li et al., 2023): A model leveraging a pre-trained GPT-2 with two-stage tuning for generating explanations. We specifically use the PEPLER-MF variant, known for its superior text quality. Among these models, PEPLER’s architecture closely resembles ours, making it a crucial point of comparison.
  • ERRA (Cheng et al., 2023): A model that combines both aspect-modeling and retrieval into the revised transformer architecture. While ERRA and MAPLE share similar core ideas, our model separates the retrieval process from itself and employs an LLM reader to comprehend the retrieved information from a broader perspective.
@@
  • PRAG (Xie et al., 2023): A transformer-based model integrates user and item embeddings and review history. It employs a personalized attention mechanism to generate a latent query, which is optimized in an auto-encoder fashion against the embedding of the ground-truth review.
@@ C.1 Implementation Details
MAPLE is trained on the training set with hyperparameters fine-tuned on the validation set. Evaluation is performed on 10,000 test-set pairs, averaged over 5 splits for MAPLE and review-generation models, while PRAG is tested only on the first split due to its long training time. MAPLE is optimized using the AdamW optimizer (Loshchilov and Hutter, 2019), with a learning rate of 0.001 and a batch size of 196. The maximum sequence length is limited to 20 Byte Pair Encoding (BPE) (Sennrich et al., 2016) tokens. The training process is divided into two stages as discussed in Section 3.5: Stage 1 runs for up to 30 epochs with the tolerance times set to 5 (monitoring only the text loss); Stage 2 lasts for a maximum of 20 epochs with the tolerance times set to 2 (monitoring only the aspect-recommendation loss). We use different training datassets for each stage. In the first stage, each user-item pair is associated with various aspect categories $c_{i}$ from the training set. In the second stage, focused on training the recommendation task, we trim the dataset so that each user-item pair appears only once. The aspect regularization coefficient $\alpha$ is set to 0.01. The embedding dimension for user ID, item ID, and aspect is set to 768. For the aspect-recommendation component, we use an MLP with two hidden layers of 256 and 128 dimensions. Regarding the Distribution Balance loss, we set the negative-tolerance coefficient $\lambda$ to 1.0 and class bias $\nu_{i}$ to 0.05 for each class.
During inference, we apply greedy decoding for all models. For MAPLE, it sets $K$ to 3 as the number of aspects to be mixed as the guiding signal. For the personalized retriever of PRAG, we employ its item-marginalization variant and train with the user, item ID dimension of 768; the rest follow the default settings. For all the other baseline models, we adopt the official implementation as well as the default hyperparameter settings.
C.2 Computational Resources and Cost Analysis
All experiments are run on a single NVIDIA GeForce RTX 3090 (24 GB) GPU.
Training and Inference Costs. It takes about 14 and 7 hours to train MAPLE on Yelp23 and Yelp19, respectively. At inference, MAPLE performs aspect recommendation and then generates explanations conditioned on the predicted aspect vector. This adds only 0.02s latency per batch of 200 samples, making it computationally comparable to baseline models (e.g., PETER, PEPLER, and ERRA) under the same backbone.
+RAG Variant Resource Usage. In case studies (Appendix D), the MAPLE +RAG variant performs retrieval by encoding all reviews using a sentence transformer and computing similarity scores over a review subset. This process requires approximately 60 GB of CPU RAM. For language generation, we use the GPT-4 model via ChatGPT (accessed on June 10, 2024) with default settings and cleared history.
@@ C.2 Computational Resources and Cost Analysis
Figure 6: Prompt template for the reader model. In terms of the personal query, for MAPLE+RAG, we use the MAPLE-generated explanation; for PRAG+RAG, we use the translated and filtered keyword set.
@@ Appendix D Qualitative Case Studies
To address the issues of under-representation of user roles, we select several examples under the same item for the case study. Since the personalized retriever in Xie et al. (2023) produces a latent query for a user-item pair, we train the embedding interpreter as Xie et al. (2023) proposes to translate the query to a set of keywords: We input the latent query into the embedding interpreter and list the output keywords. In the implementation, it generates 5 keywords and then keeps only those appearing in the retrieved reviews, therefore, the displayed keywords may be fewer than five. In the enhanced +RAG model, we treat the prefix model as the retriever and utilize GPT-4 as the reader. We augment only MAPLE and PRAG with this setup. For MAPLE, we encode its explanation with all-mpnet-base-v2  [^12] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 and then treat them as queries to fetch the top-10 most similar reviews from the user and item history review pools, respectively. See 7 for illustration. For PRAG personalized retriever, we use its proposed item-marginalization variant (i.e., subtract the mean of latent queries generated for the same item, and use it as the query embedding so as to highlight the /textitdistinctiveness of each item) to retrieve pertinent reviews from the user and item pool.
D.1 Case Study 1: Item Feature Precision
It can be observed that in Table 7, PRAG’s retriever often produces queries focused narrowly on service speed, which lacks the breadth to capture the restaurant’s primary attractions. Consequently, the PRAG+RAG mostly centers the discussions on service and wait times. Some more cases show that PRAG tends to prioritize service aspects. Conversely, MAPLE excels in identifying detailed features such as ice cream flavors (e.g., salted caramel, chocolate chips, strawberry), styles (e.g., twist), and its complement (e.g., the waffle cone), offering a variety of explanations tailored to different user preferences. In contrast, other review-generation models generally produce generic terms like "ice cream" or "toppings"; the best is probably PETER’s "soft-served" in recommending user xC-q_yh0XwcjRLimkS3RNg.
@@ D.2 Case Study 2: Personalization
In exploring the personalization capabilities of recommendation systems, we focus on how effectively restaurant features are remembered and presented in user-specific contexts, i.e., how item IDs impact the generated explanations. A pertinent question arises: How about the impacts of user IDs? We assume an imaginary baseline approach, which simply samples reviews from an item’s review history. Could it possibly be as good as MAPLE or the other review-generation models? Or let’s put it this way: what are the differences between it and the other review-generation models? We observe that user signals primarily influence the style and tone of the generated sentences. For instance, in Table 8, the user has a review history predominantly in French. Learning that the past reviews are primarily in French, the models generate sentences also in French. This level of personalization would be challenging to achieve with this random sampling baseline from the item’s review history. Could such a baseline possibly rival in terms of personalization of MAPLE or other user ID-incorporated review-generation models? Moreover, MAPLE demonstrates an ability to retain and reflect personal details. For example, in Table 9, the user frequently mentions "Phoenix" in the review history, suggesting they might reside in the City of Phoenix. MAPLE captures this detail and incorporates it into the generated explanation, highlighting the user’s local preferences. This ability to integrate personal context into recommendations underscores the model’s strength in tailoring content to individual users. These observations lead to the conclusion that the random baseline could not replicate the nuanced personalization achieved by models like MAPLE, which possess sophisticated mechanisms for personalizing generated sentences, adapting not only to users’ linguistic preferences but also to subtle personal details, thereby enhancing the relevance and effectiveness of personalized recommendations.
@@
Table 7: Case studies on item id N_2yEZ41g9zDW_gWArFiHw (Sweet Jesus, an ice-cream shop) in Yelp19. There are a total of 35 reviews in the 10,000 test samples, 4 reviews are selected for demonstration. The recommended aspect terms mentioned in the explanations are boldfaced.